OpenStudy (anonymous):

When solving a system of equations, Jared found y = x + 10 for one equation and substituted x + 10 for y in the other equation. Nicole found x = y – 10 for the same equation and substituted y – 10 for x in the other equation. Who is correct? Explain.

OpenStudy (anonymous):

They both are.

OpenStudy (anonymous):

The equation (y=x+10) = (x=y-10)